The following table shows the income distribution of 600 families. Find the minimum income


of the riches 30% families. Also the limits of income of middle 50% of families, to the nearest


rupees.


Income Below


75


75-


150


150-


225


225-


300


300-


375


375-


400


400 &


above


No. of


families


69 137 225 46 88 25 10


Ans.: the richest 30 % families earns Rs. 222 and above per week , the middle 50% families


weekly income lies between 120 and 256.



1
Expert's answer
2022-01-11T13:49:50-0500

Income f cf

0-75 69 69

75-150 137 206

150-225 225 431

225-300 46 477

300-375 88 565

375-400 25 590

400+ 10 600

"a)"

To find the minimum income of the richest 30% families, we find the 70th percentile as follows.

"P_{70}=l+({70\times n\over100}-cf)\times{c\over f}" where, "n=600"

"l" is the lower class boundary of the class containing "P_{70}"

"f" is the frequency of the class containing "P_{70}"

"c" is the width of the class containing "P_{70}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{70}"

"P_{70}" is in the "({70\times600\over100})^{th} =420^{th}" position. Therefore, it lies in the class, 150-225

Thus,

"P_{70}=150+(420-206)\times{75\over225}=150+71.33=221.33\approx 222" ( to the nearest rupees)

Therefore, the richest 30% families earns Rs. 222 and above per week

"b)"

The middle 50% families lies between the 25th and the 75th percentile. To find the range of their income, we find "P_{25}" as the lower limit and "P_{75}" as the upper limit.

Now,

"P_{25}=l+({25\times n\over100}-cf)\times{c\over f}" where, "n=600"

"l" is the lower class boundary of the class containing "P_{25}"

"f" is the frequency of the class containing "P_{25}"

"c" is the width of the class containing "P_{25}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{25}"

"P_{25}" is in the "({25\times600\over100})^{th} =150^{th}" position. Therefore, it lies in the class, 75-150

Thus,

"P_{25}=75+(150-69)\times{75\over137}=75+44.34=119.34\approx 120" ( to the nearest rupees)

and,

"P_{75}=l+({75\times n\over100}-cf)\times{c\over f}" where, "n=600"

"l" is the lower class boundary of the class containing "P_{75}"

"f" is the frequency of the class containing "P_{75}"

"c" is the width of the class containing "P_{75}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{75}"

"P_{75}" is in the "({75\times600\over100})^{th} =450^{th}" position. Therefore, it lies in the class, 225-300

Thus,

"P_{75}=225+(450-431)\times{75\over46}=225+=255.98\approx 256" ( to the nearest rupees)

Therefore, the middle 50% families weekly income lies between 120 and 256.